# Game-based Platforms for Artificial Intelligence Research

### v3

@@ V-G Roles of Games in Interdisciplinary Research in AI
VI Conclusion
This paper reviews and categorises the publicly available games and game-based platforms for AI research. Growing research interests in AI for game testing, and AI for game design, games for artificial general intelligence, games for creative AI, have been induced. We present the games and game-based platforms for playing according to their characteristics, such as aims, the number of agents, observability, programming languages, and whether AI related competitions are held. Games and platforms for design are grouped by focused content type, such as 2D levels, 3D levels, narratives, rhythm charts, and cards.

### v4

@@ 5.7 Roles of Games in Interdisciplinary Research in AI
6 Conclusion
This paper reviews and categorises the publicly available games and game-based platforms for AI research. Growing research interests in AI for game testing, and AI for game design, games for artificial general intelligence, multi-agent systems, and games for creative AI, have been induced. We present the games and game-based platforms for playing according to their characteristics, such as aims, the number of agents, observability, programming languages, and whether AI related competitions were held. Games and platforms for design are grouped by focused content type, such as 2D levels, 3D levels, narratives, rhythm charts, and cards.
@@ 6 Conclusion
Acknowledgement
Authors would like to thank all anonymous reviewers for their careful review and insightful comments.